Tc-99 has a half of 6 hours. If 400 grams of tc-99 is injected into you, how much remains when you are released from the hospital 18 hours later
1
Expert's answer
2020-12-14T14:41:09-0500

The general formula is:

"N(t)=N_0(\frac{1}{2})^{\frac{t}{t_{1/2}}}"

where N(t) is the amount at given time, N0 is the initial amount, t is the given time and t1/2 is the half-life.


Therefore,


"N(18hrs)=400g(\frac{1}{2})^{\frac{18}{6}}=50.0g"


Answer: 50.0 g